solve the equation the ages of a and b are in ratio 5:7.four years from now the ratio of their ages will be 3:4 . the present age of b is

Answer by scott8148(6628)

5b = 7a

3(b + 4) = 4(a + 4) ___ 3b + 12 = 4a + 16 ___ 21b + 84 = 28a + 112

subtracting 4 times first eqn ___ 21b + 84 - 20b = 28a + 112 - 28a ___ b = 28